OpenStudy (mirandaregina):

I need help finding the y-values of the function y=sin2(x+pi/2).

OpenStudy (mirandaregina):

Here's what I have so far: Amp = 1, Horizontal Shift = pi/2 to the left, Vertical Shirt = none, Period = 2pi/2 = pi. Then I divided the period (pi) by 4 to get pi/4, and translated the x-values of the parent graph y=sinx to pi/4, pi/2, 3pi/4, pi, and 5pi/4. How do I find the y-values from here?

OpenStudy (jdoe0001):

*cough* y-values = range *cough*

OpenStudy (mirandaregina):

@Loser66 ok. What I need help with is finding how to get the y-values shown in the image attached. I've already found the x-values, though.

OpenStudy (mirandaregina):

@jdoe0001 isn't the range -1, 1?

OpenStudy (freckles):

do you mean find the associated y-values to the x-values pi/4,pi/2,3pi/4, pi,5pi/4?

OpenStudy (mirandaregina):

@freckles yes

OpenStudy (jdoe0001):

let us think about it... for a bit what's \((-1)^2\)? or what is \((+1)^2\)? keep in mind that, sine function or sin(x) has a range of \(-1\le x \le +1\)

OpenStudy (freckles):

yo can plug them into y=sin2(x+pi/2) and does this mean y=sin^2(x+pi/2) or do you mean y=sin(2x+pi/2) or do you mean y=sin(2(x+pi/2))

OpenStudy (jdoe0001):

well... so... whatever \(sin^2\) is going to spit out, will never be greater than that, and thus, that's the range

OpenStudy (mirandaregina):

@freckles no, it's actually